# Self-Supervised Autoencoder Network for Robust Heart Rate Extraction from Noisy Photoplethysmogram: Applying Blind Source Separation to Biosignal Analysis

### v1

@@ Abstract
Biosignals can be viewed as mixtures measuring particular physiological events, and blind source separation (BSS) aims to extract underlying source signals from mixtures. This paper proposes a self-supervised multi-encoder autoencoder (MEAE) to separate heartbeat-related source signals from photoplethysmogram (PPG), enhancing heart rate (HR) detection in noisy PPG data. The MEAE is trained on PPG signals from a large open polysomnography database without any pre-processing or data selection. The trained network is then applied to a noisy PPG dataset collected during the daily activities of nine subjects. The extracted heartbeat-related source signal significantly improves HR detection as compared to the original PPG. The absence of pre-processing and the self-supervised nature of the proposed method, combined with its strong performance, highlight the potential of BSS in biosignal analysis.
@@ 2.4 Validating Noisy PPG Heart Rate Extraction
The HR values were detected in 48-second segments and then combined to form beat-by-beat HR arrays for ECG, PPG, and the source signal from the selected encoder for each recording. Median smoothing with window size of 5 cardiac cycles was applied to the detected HR in order to smooth out occasional undetected beats. To compare the extracted HR from the original PPG signal, the source signal, and the reference ECG signal, root mean squared error (RMSE) and Pearson correlation values were calculated for each of the 108 recordings. Direct comparison is only made to the results of supervised bidirectional recurrent denoising autoencoder (BRDAE) method in our previous study, and alternative methods are omitted here as the proposed method outperformed the BRDAE, which significantly outperformed other methods including bandpass filtering and wavelet denoising.
@@ 3 Results
The RMSE and the correlation values calculated between the original noisy PPG and reference ECG in the daily routine testing dataset were $14.4\pm 10.6\text{BPM}$ and $0.407\pm 0.177$, respectively. For the selected PPG source signal, the values were $4.9\pm 5.1\text{BPM}$ and $0.740\pm 0.153$. As compared to our previous PPG denoising result using BRDAE, the results showed improvement in both the correlation and RMSE values. Bland-Altman plots of the HR errors for both the original PPG and the selected source signal on the same dataset is shown in Figure 5.
@@ 4.3 Limitations
5 Data and code availability statement
The MESA dataset is available from the National Sleep Research Resource: www.sleepdata.org The noisy PPG dataset is not publicly available due to the absence of permission from the study participants and approval from the relevant Institutional Review Board (IRB). Our implementation of the experiments described in this paper is available on GitHub at the following link: github.com/webstah/meae-heart-rate-extraction-from-noisy-ppg

### v2

@@ Abstract
Biosignals can be viewed as mixtures measuring particular physiological events, and blind source separation (BSS) aims to extract underlying source signals from mixtures. This paper proposes a self-supervised multi-encoder autoencoder (MEAE) to separate heartbeat-related source signals from photoplethysmogram (PPG), enhancing heart rate (HR) detection in noisy PPG data. The MEAE is trained on PPG signals from a large open polysomnography database without any pre-processing or data selection. The trained network is then applied to a noisy PPG dataset collected during the daily activities of nine subjects. The extracted heartbeat-related source signal significantly improves HR detection as compared to the original PPG. The absence of pre-processing and the self-supervised nature of the proposed method, combined with its strong performance, highlight the potential of MEAE for BSS in biosignal analysis.
@@ 2.4 Validating Noisy PPG Heart Rate Extraction
The HR values were detected in 48-second segments and then combined to form beat-by-beat HR arrays for ECG, PPG, and the source signal from the selected encoder for each recording. Median smoothing with window size of 5 cardiac cycles was applied to the detected HR in order to smooth out occasional undetected beats. To compare the extracted HR from the original PPG signal, the source signal, and the reference ECG signal, root mean squared error (RMSE) and Pearson correlation values were calculated for each of the 108 recordings. The results from the proposed method are compared to traditional BSS methods including Independent Component Analysis (ICA) and Non-negative Matrix Factorization (NMF). As the PPG recordings are one dimensional and both the ICA and the NMF requires as many observations as components, 8 pseudo-copies were made by shifting the recording 1 sample per copy, and the best performing components for each method in terms of HR error were selected for comparison. Furthermore, comparison is only made to the results of supervised bidirectional recurrent denoising autoencoder (BRDAE) method in our previous study, and alternative methods in that study are omitted here as the proposed method outperformed the BRDAE, which significantly outperformed other methods including bandpass filtering and wavelet denoising.
@@ 3 Results
The RMSE and correlation values for heart rate (HR), calculated between the original noisy PPG and the reference ECG in the daily routine testing dataset, were $14.4\pm 10.6$ BPM and $0.407\pm 0.177$, respectively. For the selected PPG source signal, the values were $4.9\pm 5.1$ BPM and $0.740\pm 0.153$. These results, along with the results from ICA, NMF, and BRDAE, are summarized in Table 1. Bland-Altman plots of the HR errors for both the original PPG and the selected source signal on the same dataset is shown in Figure 5.
Table 1: Comparison of heart rates from various methods against reference ECG HR.
None ICA NMF BRDAE MEAE
Correlation $0.407\pm 0.177$ $0.602\pm 0.214$ $0.576\pm 0.181$ $0.686\pm 0.147$ $0.740^{*}\pm 0.153$
RMSE $14.4\pm 10.6$ $8.6\pm 10.4$ $8.2\pm 6.0$ $6.1\pm 8.2$ $4.9^{*}\pm 5.1$
*Significant improvement as compared to the original PPG (p<0.01).
@@ 4.3 Limitations
5 Code availability statement
Our implementation of the experiments described in this paper is available on GitHub at the following link: github.com/webstah/meae-heart-rate-extraction-from-noisy-ppg
@@ 5 Code availability statement
6 Ethics statement
MESA dataset
The MESA dataset is available from the National Sleep Research (www.sleepdata.org) following an application process for access.
9-subject noisy photoplethysmogram dataset
This study was conducted in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki declaration and its later amendments or comparable ethical standards. All subjects provided informed consent and the study was approved by the Institutional Review Board of Seoul National University Hospital.